# thanks: Corresponding authors:

### v1

@@ A.1 Coarse-graining the auto-activation model
The statistical mechanical model for autoactivation depicted in Fig. 2 implicitly accounts for interaction between the activator and polymerase. Fig. 26 represents the complete accounting of the thermodynamic states, weights and rates, now explicitly accounting for all of the possible polymerase bound states and denoting the interaction energy between polymerase and activator as $\varepsilon_{ap}$. In light of this complete set of states, weights and rates, We can write the dynamical equation for the number of activators as
@@ B.2 Necessary condition for bistability for a fixed concentration of effector $c$
$\begin{cases}c_{\text{bistab}}^{\max}>c>c_{\text{bistab}}^{\min}&\text{if }\frac{1}{1+e^{-\beta\varepsilon}\bar{K}_{c}^{2}}<\frac{2}{\omega\bar{r}_{2}}<\frac{1}{2\bar{r}_{1}}<\frac{1}{1+e^{-\beta\varepsilon}},\\ c_{\text{bistab}}^{\max}>c&\text{if }\frac{1}{1+e^{-\beta\varepsilon}\bar{K}_{c}^{2}}<\frac{2}{\omega\bar{r}_{2}}<\frac{1}{1+e^{-\beta\varepsilon}}<\frac{1}{2\bar{r}_{1}},\\ c>c_{\text{bistab}}^{\min}&\text{if }\frac{2}{\omega\bar{r}_{2}}<\frac{1}{1+e^{-\beta\varepsilon}\bar{K}_{c}^{2}}<\frac{1}{2\bar{r}_{1}}<\frac{1}{1+e^{-\beta\varepsilon}},\\ c\geq 0&\text{if }\frac{2}{\omega\bar{r}_{2}}<\frac{1}{1+e^{-\beta\varepsilon}\bar{K}_{c}^{2}}<\frac{1}{1+e^{-\beta\varepsilon}}<\frac{1}{2\bar{r}_{1}},\\ \text{no bistability}&\text{if }\frac{2}{\omega\bar{r}_{2}}>\frac{1}{1+e^{-\beta\varepsilon}}\text{ or }\frac{1}{1+e^{-\beta\varepsilon}\bar{K}_{c}^{2}}>\frac{1}{2\bar{r}_{1}}.\end{cases}$ (89)
@@ Appendix C Fixed point structure of the auto-activation system as a gradient flow
$-\frac{dV}{d\bar{A}}=\frac{P(\bar{A})}{1+2p_{\text{act}}(c)\bar{A}+\omega(p_{\text{act}}(c)\bar{A})^{2}},$ (97)
@@
$-dV=d\bar{A}\cdot\frac{1+2p_{\text{act}}(c)\bar{A}+\omega(p_{\text{act}}(c)\bar{A})^{2}}{\omega p_{\text{act}}^{2}(c)(\bar{A}-\bar{A}_{1})(\bar{A}-\bar{A}_{2})(\bar{A}-\bar{A}_{3})}.$ (99)
@@
Figure 29: Parameter space exploration of the maximal cooperativity $\omega_{\max}^{\text{bistable}}$ above which the system becomes monostable for all effector concentrations. The cooperativity $\omega$ is sampled over the interval $\omega\in[10^{1},10^{9}]$. Effector concentrations are effectively scanned from 0 to $\infty$ by sweeping $p_{\text{act}}$ between its biologically constrained bounds: the leakiness level $p_{\text{act}}^{\text{min}}=\frac{1}{1+e^{-\beta\varepsilon}\bar{K}_{c}^{2}}$ and the saturation level $p_{\text{act}}^{\text{max}}=\frac{1}{1+e^{-\beta\varepsilon}}$, with fixed parameters $\beta\varepsilon=4.5$ and $\bar{K}_{c}=2.6\times 10^{2}$. Three two-dimensional parameter sweeps are performed. In panel (A), $(\bar{r}_{1},\bar{r}_{2})$ are varied in $[\bar{r}_{0},10^{5}]\times[\bar{r}_{0},10^{5}]$ with $\bar{r}_{0}=0.1$ held constant. In panel (B), $(\bar{r}_{0},\bar{r}_{2})$ are varied in $[10^{-5},\bar{r}_{1}]\times[\bar{r}_{1},10^{5}]$ with $\bar{r}_{1}=1$ fixed. In panel (C), $(\bar{r}_{0},\bar{r}_{1})$ are varied in $[10^{-5},\bar{r}_{2}]\times[10^{-5},\bar{r}_{2}]$ with $\bar{r}_{2}=20$ fixed. Regions shaded in gray correspond to parameter combinations that violate the auto-activation condition $\bar{r}_{0}\leq\bar{r}_{1}\leq\bar{r}_{2}$, and for which the system no longer functions as an auto-activating unit. In regions where no maximal cooperativity values for bistability are reported, the system remains monostable across the entire range of cooperativity values sampled.
@@
We now study the asymptotic behavior of this discriminant in the limits of infinite $\omega$, $\bar{r}_{2}$, and $\bar{r}_{0}$. We do not consider the limit of infinite $\bar{r}_{1}$, since, according to the bounds in Eqns. 1012, a necessary condition for bistability is that $\bar{r}_{1}$ remains below a threshold set by the other system parameters. Therefore, in this limit, the system is necessarily monostable.
@@ Appendix E Conditions for activation in auto-activation circuit
Appendix F Auto-activation: Supplement for the comparison between Hill and thermodynamic model
Figure 30: Comparing probabilities of activator binding states as a function of active activator concentration $\bar{A}_{\text{act}}$ in the thermodynamic and Hill model. (A) Thermodynamic model with cooperativity $\omega=1$. (B) Thermodynamic model with $\omega=100$. (C) Hill function model with Hill coefficient $n=2$. Blue curve corresponds to the state with no activator bound. Orange curve corresponds to the state with one activator bound. Green curve corresponds to the state with two activators bound. Notably, there is no orange curve in (C) as the Hill function approximates away the state with one activator bound.
@@ Appendix F Auto-activation: Supplement for the comparison between Hill and thermodynamic model
In this section, we provide support for some of the claims made in the discussion comparing thermodynamic models and the Hill function approach in the context of auto-activation in Sec. III.1.2.
First, we address the claim that the Hill function formulation produces similar probabilities of state of transcription factor binding. In Fig. 30, we illustrate such probabilities as a function of active activator concentration $\bar{A}_{\text{act}}=p_{\text{act}}(c)\bar{A}$. For example, the probability of no activator bound in the thermodynamic model is given by
$\frac{1}{1+2\bar{A}_{act}+\omega(\bar{A}_{act})^{2}}.$ (114)
Here, we treat the Hill function as approximating away the state where one activator is bound, and only allows zero or two activators to be bound. We observe that the probability of no activator bound (blue curve) and the probability of two activators bound (green curve) have similar shapes in both models. The two models differ slightly in the region where the green and blue curves intersect which signifies the transition region of the genetic switch. In the thermodynamic model, the green and blue curves typically intersect at a lower probability due to the existence of the state where one activator is bound, and the $\bar{A}_{\text{act}}$ at which they intersect can be tuned by $\omega$. Nevertheless, it is difficult to intuit the drastic differences in bifurcation curves between the two models as described in the main text from this picture, as the qualitative features of the probabilities of states in the two models resemble each other.
Figure 31: Hill function predicts similar EC50 in their bifurcation curves for all Hill coefficients $n$. Parameters used here are the same as in Fig. 12(B), with $\bar{r}_{0}=0.1$ and $\bar{r}_{2}=10$. Bifurcation curves with different Hill coefficients are colored differently; blue curve: $n=2$; orange curve: $n=5$; green curve: $n=10$. It can be observed that the critical effector concentration $c$ at which the activator concentration transitions from high to low steady state are similar for all three bifurcation curves.
Second, in the main text (Fig. 12(B)), we illustrated the behavior of the Hill function formulation using a representative example with Hill coefficient $n=2$. To support the generality of the observation that the critical effector concentration $c$—captured by the half-maximal effective concentration EC50—remains approximately fixed across Hill coefficients, we include in Fig. 31 additional bifurcation curves for $n=2,5,10$. These curves demonstrate that EC50 varies little with $n$ for Hill curves. By contrast, in the thermodynamic model, the location of EC50 is sensitive to parameters such as $\bar{r}_{1}$ and $\omega$, as shown in Fig. 12(B).
@@
Appendix G Auto-activation : Bistability is possible for non cooperative systems ($\omega=1$).
G.1 Definition of the effective Hill coefficient

### v2

@@ A.1 Coarse-graining the auto-activation model
The statistical mechanical model for auto-activation depicted in Fig. 2 implicitly accounts for interaction between the activator and polymerase. Fig. 26 represents the complete accounting of the thermodynamic states, weights and rates, now explicitly accounting for all of the possible polymerase bound states and denoting the interaction energy between polymerase and activator as $\varepsilon_{ap}$. In light of this complete set of states, weights and rates, We can write the dynamical equation for the number of activators as
@@ B.2 Necessary condition for bistability for a fixed concentration of effector $c$
[…]
@@ Appendix C Fixed point structure of the auto-activation system as a gradient flow
$\frac{dV}{d\bar{A}}=\frac{P(\bar{A})}{1+2p_{\text{act}}(c)\bar{A}+\omega(p_{\text{act}}(c)\bar{A})^{2}},$ (97)
@@
$dV=d\bar{A}\cdot\frac{1+2p_{\text{act}}(c)\bar{A}+\omega(p_{\text{act}}(c)\bar{A})^{2}}{\omega p_{\text{act}}^{2}(c)(\bar{A}-\bar{A}_{1})(\bar{A}-\bar{A}_{2})(\bar{A}-\bar{A}_{3})}.$ (99)
@@
Figure 29: Parameter space exploration of the maximal cooperativity $\omega_{\max}^{\text{bistable}}$ above which the system becomes monostable for all effector concentrations. The cooperativity $\omega$ is sampled over the interval $\omega\in[10^{1},10^{9}]$. Effector concentrations are effectively scanned from 0 to $\infty$ by sweeping $p_{\text{act}}$ between its biologically constrained bounds: the leakiness level $p_{\text{act}}^{\text{min}}=\frac{1}{1+e^{-\beta\varepsilon}\bar{K}_{c}^{2}}$ and the saturation level $p_{\text{act}}^{\text{max}}=\frac{1}{1+e^{-\beta\varepsilon}}$, with fixed parameters $\beta\varepsilon=4.5$ and $\bar{K}_{c}=2.6\times 10^{2}$. Three two-dimensional parameter sweeps are performed. In panel (A), $(\bar{r}_{1},\bar{r}_{2})$ are varied in $\bar{r}_{0},10^{5}]\times[\bar{r}_{0},10^{5}]$ with $\bar{r}_{0}=0.1$ held constant. In panel (B), $(\bar{r}_{0},\bar{r}_{2})$ are varied in $10^{-5},\bar{r}_{1}]\times[\bar{r}_{1},10^{5}]$ with $\bar{r}_{1}=1$ fixed. In panel (C), $(\bar{r}_{0},\bar{r}_{1})$ are varied in $10^{-5},\bar{r}_{2}]\times[10^{-5},\bar{r}_{2}]$ with $\bar{r}_{2}=20$ fixed. Regions shaded in gray correspond to parameter combinations that violate the auto-activation condition $\bar{r}_{0}\leq\bar{r}_{1}\leq\bar{r}_{2}$, and for which the system no longer functions as an auto-activating unit. In regions where no maximal cooperativity values for bistability are reported, the system remains monostable across the entire range of cooperativity values sampled.
@@
We now study the asymptotic behavior of this discriminant in the limits of infinite $\omega$, $\bar{r}_{2}$, and $\bar{r}_{0}$. We do not consider the limit of infinite $\bar{r}_{1}$, since, according to the bounds in Eqns. 10 - 12, a necessary condition for bistability is that $\bar{r}_{1}$ remains below a threshold set by the other system parameters. Therefore, in this limit, the system is necessarily monostable.
@@ Appendix E Conditions for activation in auto-activation circuit
Appendix F Comparison between the Hill and thermodynamic models for auto-activation
[…]
@@ Appendix F Comparison between the Hill and thermodynamic models for auto-activation
In this section, we provide support for some of the claims made in Section III.1.2 when comparing the use of thermodynamic and Hill function approaches to model auto-activation.
First, we demonstrate how a Hill function emerges in the high cooperativity limit of the thermodynamic model, a result known previously in the literature [90]. To reiterate, the dynamical equation for protein production through auto-activation using a thermodynamical model is
$\begin{split}\frac{dA}{dt}=-\gamma A&+\frac{r_{0}+r_{1}(2p_{\text{act}}(c)\frac{A}{K_{d}})+r_{2}\omega(p_{\text{act}}(c)\frac{A}{K_{d}})^{2}}{1+2p_{\text{act}}(c)\frac{A}{K_{d}}+\omega(p_{\text{act}}(c)\frac{A}{K_{d}})^{2}}.\end{split}$ (114)
Letting $K_{d}^{\text{eff}}=K_{d}/\sqrt{\omega}$, we can then re-express the previous equation as
$\frac{dA}{dt}=-\gamma A+\frac{r_{0}+r_{1}2p_{\text{act}}(c)\frac{A}{K_{d}^{\text{eff}}\sqrt{\omega}}+r_{2}(p_{\text{act}}(c)\frac{A}{K_{d}^{\text{eff}}})^{2}}{1+2p_{\text{act}}(c)\frac{A}{K_{d}^{\text{eff}}\sqrt{\omega}}+(p_{\text{act}}(c)\frac{A}{K_{d}^{\text{eff}}})^{2}}.$ (115)
In the limit as $\omega\to\infty$ with finite $K_{d}^{\text{eff}}$, the single activator-bound state vanishes and the dynamics simplify to
$\frac{dA}{dt}=-\gamma A$ $+\frac{r_{0}+r_{2}(p_{\text{act}}(c)\frac{A}{K_{d}^{\text{eff}}})^{2}}{1+(p_{\text{act}}(c)\frac{A}{K_{d}^{\text{eff}}})^{2}}.$ (116)
The production term thus takes the form of a Hill function with $n=2$, as in Eqn. 14.
We next compare the probabilities of different state occupancies as derived from the thermodynamic and Hill function models. Fig. 30 plots the probabilities of no TF bound, one TF bound, and two TFs bound as functions of the active activator concentration $A_{\text{act}}=p_{\text{act}}(c)A$. In the thermodynamic model the probability of no activator bound is
$\frac{1}{1+2\frac{A_{\text{act}}}{K_{d}}+\omega\left(\frac{A_{\text{act}}}{K_{d}}\right)^{2}}.$ (117)
[…]
Overall, these differences appear moderate, and it is difficult to anticipate from these results alone the substantive differences between the two models in the small $\omega$ regime that we present in the main text. This highlights the importance of comparing the two models not only through their production curves but also through their bifurcation curves.
Appendix G Oligomerization and DNA Looping in the auto-activation regulation unit
In the main text, transcription factors (TFs) are treated as single binding units that occupy one site at a time, with cooperative interactions between TFs bound at two neighboring sites on the DNA. This description, however, omits two important possibilities: (i) TFs may oligomerize before binding DNA (for example as dimers, like the transcription factor PhoP in E. coli [72], or as tetramers, like the transcription factor ComK [35] in B. subtilis); and (ii) they may bridge distant sites to form DNA loops [64, 3, 99, 100]. Here we consider how these features would change the behavior of an auto-activating transcription factor, with similar analysis possible for mutual repression and other regulatory motifs.
To isolate the effects of oligomerization, the following discussion does not include inducers, though in general they could interact differently with monomers and oligomers. To extend the model, Eqn. 8 in Section III.1, we now allow the transcription factor to dimerize, and each protein copy to exist as a free monomer or as part of a dimer. We track the total concentration of monomer subunits as
$A_{\text{tot}}=A+2A_{2},$ (118)
where $A_{2}$ denotes dimers. Protein synthesis increases $A_{\text{tot}}$, and assuming rapid dimerization the newly produced monomers are instantly redistributed between monomers and dimers according to equilibrium. We assume that dilution and degradation act at the same effective rate $\gamma$ on both monomers and dimers, which is reasonable for stable TFs in fast-growing E. coli where dilution dominates, as a few candidates of auto-activating feedback loops are present in this organism [33, 86].
Assuming the reaction $2A\leftrightarrow A_{2}$ is much faster than synthesis or degradation, the dimer concentration is given by
$A_{2}=\frac{A^{2}}{K_{\text{dim}}}.$ (119)
Mass balance then gives
$A_{\text{tot}}=A+\frac{2A^{2}}{K_{\text{dim}}},$ (120)
which, solving for $A$, yields an explicit definition for the free monomer concentration as a function of $A_{\text{tot}}$,
$A(A_{\text{tot}})=\frac{K_{\text{dim}}}{4}\!\left(\sqrt{1+\frac{8A_{\text{tot}}}{K_{\text{dim}}}}-1\right),$ (121)
and, given Eqn. 122, the corresponding free dimer concentration as a function of $A_{\text{tot}}$,
$A_{2}(A_{\text{tot}})=\frac{A(A_{\text{tot}})^{2}}{K_{\text{dim}}}.$ (122)
Figure 31: Promoter states for a dimerizing auto-activating transcription factor. Each row shows a possible promoter occupancy state when two binding sites are available. States differ by the number and type of bound transcription factor (monomer or dimer) and by whether the two dimers form a DNA loop. Each state is associated with a statistical weight determined by its binding configuration and a transcription rate ($r_{0}$, $r_{1}$, or $r_{2}$) that depends on the number of bound molecules.
The transcription factor can bind DNA either as a monomer, with dissociation constant $K_{d}^{m}$, or as a dimer, with dissociation constant $K_{d}^{d}$. We now consider a promoter with two identical binding sites. The binding of two proteins can be cooperative, described by cooperativity parameters $\omega_{mm}$ (monomer–monomer), $\omega_{md}$ (monomer–dimer), and $\omega_{dd}$ (dimer–dimer). When both sites are occupied by dimers, they can also bridge the two sites to form a DNA loop, with a dissociation constant $K^{l}_{d}$ describing the dimer bound to DNA in its looped state. Each promoter state is associated with a transcription rate $r_{0}$, $r_{1}$, or $r_{2}$, corresponding to zero, one, or two bound proteins, respectively. Given the statistical weights of all promoter states as presented in Fig. 31, the mean transcription rate is then
$F(A,A_{2})$ $=\frac{r_{0}}{Z}+r_{1}\frac{2\frac{A}{K^{m}_{d}}+2\frac{A_{2}}{K^{d}_{d}}}{Z}$
$\quad+r_{2}\frac{\omega_{mm}\!\left(\frac{A}{K^{m}_{d}}\right)^{2}+2\omega_{md}\frac{A}{K^{m}_{d}}\frac{A_{2}}{K^{d}_{d}}+\omega_{dd}\!\left(\frac{A_{2}}{K^{d}_{d}}\right)^{2}}{Z}$
$\quad+r_{2}\frac{\frac{A_{2}}{K^{l}_{d}}}{Z},$ (123)
with a partition function
$Z=$ $1+2\frac{A}{K^{m}_{d}}+2\frac{A_{2}}{K^{d}_{d}}+\omega_{mm}\!\left(\frac{A}{K^{m}_{d}}\right)^{2}+2\omega_{md}\frac{A}{K^{m}_{d}}\frac{A_{2}}{K^{d}_{d}}$ (124)
$\quad+\omega_{dd}\!\left(\frac{A_{2}}{K^{d}_{d}}\right)^{2}+\frac{A_{2}}{K^{l}_{d}},$
and the corresponding dynamics of the system become
$\frac{dA_{\text{tot}}}{dt}=-\gamma A_{\text{tot}}+F\!\big(A(A_{\text{tot}}),A_{2}(A_{\text{tot}})\big).$ (125)
Eqn. 125 shows that once oligomerization is included, the system must be described in terms of $A_{\text{tot}}$ rather than the free monomer alone. Importantly, the dynamics become more nonlinear, allowing new behaviors to emerge.
We now consider an alternative setting in which several oligomeric forms coexist in solution, but only one can bind to DNA. As an example, suppose that monomers and dimers are found in solution with equilibrium constant $K_{\text{dim}}$, but only monomers can bind to the gene promoter site. We then have
$\frac{dA_{\text{tot}}}{dt}=-\gamma A_{\text{tot}}+\frac{r_{0}+2r_{1}\frac{A}{K^{m}_{d}}+r_{2}\omega_{mm}(\frac{A}{K^{m}_{d}})^{2}}{1+2\frac{A}{K^{m}_{d}}+\omega_{mm}(\frac{A}{K^{m}_{d}})^{2}}.$ (126)
Substituting $A_{\text{tot}}=A+2A^{2}/K_{\text{dim}}$ and using the chain rule, we can ultimately arrive at an ordinary differential equation with additional terms beyond those of the simple monomer case,
$\begin{split}\frac{dA}{dt}=&-\gamma\frac{K_{\text{dim}}A+2A^{2}}{K_{\text{dim}}+4A}\\ &+\frac{K_{\text{dim}}}{K_{\text{dim}}+4A}\frac{r_{0}+2r_{1}\frac{A}{K^{m}_{d}}+r_{2}\omega_{mm}(\frac{A}{K^{m}_{d}})^{2}}{1+2\frac{A}{K^{m}_{d}}+\omega_{mm}(\frac{A}{K^{m}_{d}})^{2}}.\end{split}$ (127)
The presence of multiple oligomeric states could therefore alter the number and stability of steady states, although in practice usually only one oligomeric form is dominant.
We next consider looping when only the dimeric form of the transcription factor is present. If a dimer both binds and loops DNA, looping can be seen as adding binding configurations that strengthen effective binding and cooperativity. The production term of a looping dimer $A_{2}$ can then be written as
$\frac{r_{0}+r_{1}(2+\frac{K^{d}_{d}}{K^{l}_{d}})p_{\text{act}}\frac{A_{2}}{K^{d}_{d}}+r_{2}\omega_{dd}\left(\frac{p_{\text{act}}A_{2}}{K^{d}_{d}}\right)^{2}}{1+(2+\frac{K^{d}_{d}}{K^{l}_{d}})p_{\text{act}}\frac{A_{2}}{K^{d}_{d}}+\omega_{dd}\left(\frac{p_{\text{act}}A_{2}}{K^{d}_{d}}\right)^{2}},$ (128)
which is algebraically equivalent to the standard model using renormalized parameters
$\tilde{K}^{d}_{d}=\frac{2K^{d}_{d}}{2+\frac{K^{d}_{d}}{K^{l}_{d}}}$ (129)
and
$\tilde{\omega_{dd}}=\frac{4\omega_{dd}}{(2+\frac{K^{d}_{d}}{K^{l}_{d}})^{2}}.$ (130)
Looping therefore does not change the functional form of the model but simply rescales $K^{d}_{d}$ and $\omega_{dd}$.
Finally, when dimers form from monomers with equilibrium constant $K_{\text{dim}}$, and there is no looping and all cooperativity terms are equal ($\omega_{mm}=\omega_{md}=\omega_{dd}=\omega$), the production term can be reduced exactly to the monomer-only functional form as
$f_{\text{mono}}(x)=\frac{r_{0}+2r_{1}x+r_{2}\,\omega x^{2}}{1+2x+\omega x^{2}},$ (131)
where
$x=\frac{A}{K^{m}_{d}}+\frac{A_{2}}{K^{d}_{d}}.$ (132)
The dynamical equation can then be written as
$\frac{dA_{\text{tot}}}{dt}=-\gamma\,A_{\text{tot}}+f_{\text{mono}}\!\big(x(A_{\text{tot}})\big).$ (133)
At steady state this means that
$\gamma(x)x=f_{\text{mono}}(x),$ (134)
with
$\gamma(x)=\frac{2K^{d}_{d}\left(K^{d}_{d}-K^{m}_{d}\left(K_{dim}+\sqrt{K^{d}_{d}\left(\frac{K^{d}_{d}}{(K^{m}_{d})^{2}}+4x\right)}\right)\right)}{K^{d}_{d}\left(1-2K_{dim}\right)-K^{m}_{d}\sqrt{K^{d}_{d}\left(\frac{K^{d}_{d}}{(K^{m}_{d})^{2}}+4x\right)}}.$ (135)
Thus, the production curve retains the same shape as in the monomeric model, but the effective degradation rate becomes concentration-dependent.
The various settings discussed in this appendix demonstrate how incorporating oligomerization alters the equations of the dynamical system. Specifically, accounting for oligomerization introduces additional molecular states, and makes the effective production term a function of the total protein pool rather than only the active monomer. This transforms the dynamics from a simple one-variable system into a more nonlinear one. By contrast, looping does not change the mathematical structure of the model but can be absorbed into effective parameters that strengthen binding and cooperativity. When only a single oligomeric form binds DNA, the production term can still be written in the same functional form as the monomer-only model, but the degradation term becomes effectively concentration-dependent.
@@ Appendix G Oligomerization and DNA Looping in the auto-activation regulation unit
Appendix H Auto-activation : Bistability is possible for non cooperative systems ($\omega=1$).
H.1 Definition of the effective Hill coefficient